# {NoHyper}footnotetext: *Co-first author

### v1

@@ Appendix C Further Numeric-Head Details
Our numeric head is composed of a tanh layer, followed by a linear layer, a GELU activation, and a final linear projection. The final LLaMA hidden state is passed through an RMS norm (Zhang & Sennrich, 2019) before it is shared between the token head and numeric head, which re-scales but does not re-center the embedding.
During training, the locations of these tokens in the ground-truth sequence are known so they can be masked to apply the MSE loss. During sampling, the position of these tokens is not pre-known and dependent on the generated sequence. We first generate the token-only sequence and then substitute the estimated numbers back in with a second pass.

### v2

@@ Appendix C Further Numeric-Head Details
Our numeric head is composed of a tanh layer, followed by a linear layer, a GELU activation, and a final linear projection. The final LLaMA hidden state is passed through an RMS norm (Zhang & Sennrich, 2019) before it is shared between the token head and numeric head, which rescales but does not recenter the embedding.
During training, the locations of these tokens in the ground-truth sequence are known so they can be masked to apply the MSE loss. During sampling, the position of these tokens is not pre-known and is dependent on the generated sequence. We first generate the token-only sequence and then substitute the estimated numbers back in a single additional pass.
Appendix D ShapeNet 6-DoF YOLOX-6D-Pose
While our investigation focused on the ability of our framework to generalize across distribution shifts, and our intention was not to suggest it as optimal over task-specific approaches to these particular problems, we additionally train and evaluate YOLOX-6D-Pose (Maji et al., 2024) on our ShapeNet scene-level 6-DoF setting to serve as a quantitative point of comparison against a recent state-of-the-art specialized method. YOLOX-6D-Pose is a modular approach trained to detect objects and regress their 6-DoF pose. In contrast to our framework, which learns to disentangle images into compositional scene representations solely by modeling the token sequence, YOLOX-6D-Pose relies on bounding-box supervision, known camera parameters, access to 3D assets, hand-crafted losses, and extensive task-specific data augmentation. By foregoing such inductive biases in the training of our framework, we sought to keep it deliberately generic to maintain the focus of our investigation on generalization without relying on task-specific designs, though such supervision and data augmentation could also be applied to our approach.
Following Maji et al. (2024), we train the extra-large YOLOX-6D-Pose variant for 300 epochs, enabling all components. However, the model cannot be compared directly to our framework. Because the detector scores object localizations with a confidence rather than producing a definite object set, we select the top-n scoring detections equal to the ground-truth number of objects in the scene after non-maximum suppression is performed. The method is additionally unable to compositionally estimate object attributes (i.e., color). The goal of the OOD ShapeNet evaluation was to examine the ability of our framework to generalize across visual domains with OOD texture and shape serving as a quantifiable proxy for real-world visual generalization. In contrast to YOLOX-6D-Pose, which is trained to be color-invariant through visual augmentation, our method, tasked to estimate color, must be color-variant. This complicates the evaluation of generalization ability.
Evaluation metrics are shown in Tab. S.1. YOLOX-6D-Pose excels in-distribution, estimating 6D rotations to, on average, one degree of error. However, despite its strong augmentation, this performance does not transfer out-of-distribution, with its scene-level chamfer distance increasing 19-fold in the OOD-T setting as opposed to the 2x increase in the float-based IG-LLM variant, which also outperforms it. This suggests that extensive data augmentation alone may not be enough to ensure out-of-distribution generalization.
Table S.1: ShapeNet 6-DoF YOLOX-6D-Pose Results. (Appendix D) We observe that while YOLOX-6D-Pose (Maji et al., 2024) excels in the ShapeNet setting in-distribution, it does not as consistently generalize.
ID OOD-T OOD-T+S
Maji et al. Float Maji et al. Float Maji et al. Float
[…]